OpenStudy (howard-wolowitz):

If a person rolls a sum of 8 when she rolls two dice, she wins $10. If the game is to be fair, how much (to the nearest penny) should the person pay to play the game?

OpenStudy (zarkon):

Let c be the cost of the game. then the expected winnings is (10-c)*P(roll an 8)+(-c)P(you don't get an 8)

OpenStudy (howard-wolowitz):

is there some way to simplify it down?

OpenStudy (howard-wolowitz):

what if i did this? Is this the same? 10cP+8-c

OpenStudy (zarkon):

(10-c)*P(roll an 8)+(-c)P(you don't get an 8) =10*P(roll and 8)-c set equal to zero (fair game) then c=10*P(roll an 8)

OpenStudy (howard-wolowitz):

so wouldnt the answer be $1.39

OpenStudy (phi):

yes, that looks good. to review the amount you get if you roll an 8 is 10-cost the amount you "get" is -cost if you roll a "not 8" the expected "earnings" is (10-cost)* P(8) + -cost*P(not 8) = 0 (which means in the long run, you don't win or lost any money) distribute: 10* P(8) - cost*P(8) - cost* P(not 8) = 0 factor out -cost: 10* P(8) - cost * (P(8) + P(not 8) ) =0 notice the probability of getting an 8 plus the probability of not getting 8 is 1 (100% one or then other happens) so the equation simplifies to 10*P(8) - cost *1 = 0 or cost = 10*P(8) you now need to find the probability of getting an 8 there are 5 ways (use a table to find the combinations) out of 36 different rolls. P(8)= 5/36 cost = 10*5/36